OpenStudy (anonymous):

Prove that d/dx arcsecu= 1/(usqrt(u^2-1)) du/dx

OpenStudy (anonymous):

\[\frac{ d }{ dx } arcsec (u)=\frac{ 1 }{ u \sqrt{u^{2}-1} } \frac{ du }{ dx }\]

OpenStudy (anonymous):

y=\[\sec^{-1}x\] \[secy=x\] \[\frac{ d }{ dx } secy = \frac{ d }{ dx } x\] \[(secytany) \frac{ du }{ dx } = 1\]

OpenStudy (anonymous):

The du/dx on the LHS should be dy/dx, but other than that, good so far.

OpenStudy (anonymous):

\[\frac{ dy }{ dx }= \frac{ 1 }{ secytany }\] \[\sec^{2}y-1 = \tan^{2}y -> \sqrt{\sec^{2}y-1}=tany\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ 1 }{ (secy)\sqrt{\sec^{2}y-1} } -> \frac{ 1 }{ xsqrt{x^{2}-1} } dx\]

OpenStudy (anonymous):

Is that right????

OpenStudy (anonymous):

oh I wrote sqrt thing in wrong lol

OpenStudy (anonymous):

\[\frac{ 1 }{ x \sqrt{x^{2}-1} }\]

OpenStudy (anonymous):

how do I get rid of the x

OpenStudy (anonymous):

In your second comment, you let y = arcsec(x). In fact, it should be y = arcsec(u). Just make the switch and it all works out.